OpenStudy (anonymous):

Find the distance between the given points. Then find the midpoint of the line segment whose endpoints are the giver points. (5 , -3/2), (-2 , 9/2)

OpenStudy (saifoo.khan):

use the distance formula.

OpenStudy (saifoo.khan):

http://0.tqn.com/d/math/1/0/A/1/distance.gif

OpenStudy (ybarrap):

X midpoint = xmid = average(5,-2)= (5 - 2)/2 = 3/2 Y midpoint = ymid = average(-3/2,9/2) = (-3/2 + 9/2)/2 = 6/2 / 2 = 3 /2 Ans. Midpoint (xmid,ymid) = (3/2,3/2)